4.  Problem 16.04

PRO FORMA INCOME STATEMENT

Austin Grocers recently reported the following 2016 income statement (in millions of dollars):

Sales $700
Operating costs including depreciation 500
EBIT $200
Interest 40
EBT $160
Taxes (40%) 64
Net income $96
Dividends $32
Addition to retained earnings $64

For the coming year, the company is forecasting a 25% increase in sales, and it expects that its year-end operating costs, including depreciation, will equal 65% of sales. Austin's tax rate, interest expense, and dividend payout ratio are all expected to remain constant.

  1. What is Austin's projected 2017 net income? Enter your answer in millions. For example, an answer of $13,000,000 should be entered as 13. Round your answer to two decimal places.
    $ ______million

  2. What is the expected growth rate in Austin's dividends? Do not round your intermediate calculations. Round your answer to two decimal places.
    _________%
Answer #1

Dividend payout ratio in 2016 = Dividends/ Net Income = 32/96 = 1/3

Sales in 2017 = 700 * 1.25 = $ 875

Operating costs including depreciation = 65% * 875 = 568.75

EBIT = 875 - 568.75 = $ 306.25

Interest = 40

EBT = 306.25 - 40 = $ 266.25

Tax rate = 40%

a)

Net Income in 2017 = $ 159.75

b)

Dividends in 2017 = 1/3 * 159.75 = $ 53.25

Dividend growth rate in Austin's dividends = (53.25 - 32)/ 32 * 100 = 66.41 %